OpenStudy (anonymous):

What is the sum of a 7 term geometric series if the first term is -11 and the last term is -45,056 and common ratio -4

OpenStudy (campbell_st):

well if you know the 1st term and common ratio you can use the formula \[a_{n} = a_{1} \times r^{n - 1}\] you know a1 = -11 and r = -4 and n = 7 so substitute and calculate away

OpenStudy (whpalmer4):

To get the sum of the first \(n\) terms, \[S_n =a_1\frac{(1-r^n)}{(1-r)}\]